OpenStudy (dqswag):

What is the Value of k(-6)

OpenStudy (anonymous):

k x -6

OpenStudy (anonymous):

You multiply

OpenStudy (dqswag):

k(x)= x+14 if < - 7 21 if -7 <_ x < -6 x^2 -1 if x>_ -6

OpenStudy (anonymous):

When you multiply a positive and a negative your always going to get a negative if it's a negative and negative it's always a positive.

OpenStudy (anonymous):

Just solve for x

OpenStudy (dqswag):

How??

OpenStudy (welshfella):

if x = -6 then k(x) = x^2 - 1 so you just plug in x = -6 into x^2 - 1 and work it out.

OpenStudy (welshfella):

that gives you k(-6)

OpenStudy (dqswag):

so I have to go through each promblem working them out,,, k(x)= x+14 if < - 7 21 if -7 x≤-6 x^2 -1 if x≥-6

OpenStudy (welshfella):

no this is a step wise function as you see the last line says that k(x) = x^2 - 1 when x is equal to or greater than -6 Here x = -6 so you only need to plug -6 into x^2 - 1.

OpenStudy (dqswag):

35????

OpenStudy (welshfella):

thats correct

OpenStudy (dqswag):

oh ok